# VRoPE: Rotary Position Embedding for Video Large Language Models

### v2

@@ Abstract
Rotary Position Embedding (RoPE) has shown strong performance in text-based Large Language Models (LLMs), but extending it to video remains a challenge due to the intricate spatiotemporal structure of video frames. Existing adaptations, such as RoPE-3D, attempt to encode spatial and temporal dimensions separately but suffer from two major limitations: positional bias in attention distribution and disruptions in video-text transitions. To overcome these issues, we propose Video Rotary Position Embedding (VRoPE), a novel positional encoding method tailored for Video-LLMs. Specifically, we introduce a more balanced encoding strategy that mitigates attention biases, ensuring a more uniform distribution of spatial focus. Additionally, our approach restructures positional indices to ensure a smooth transition between video and text tokens. Extensive experiments on different models demonstrate that VRoPE consistently outperforms previous RoPE variants, achieving significant improvements in video understanding, temporal reasoning, and retrieval tasks. Code will be available at https://github.com/johncaged/VRoPE.
@@ 1 Introduction
To optimally encode positional relationships in Video-LLMs, we identify three key properties that an ideal video positional encoding should satisfy:
@@
(1) Spatiotemporal Structure Modeling. Unlike text, where positional relationships are strictly one-dimensional, video frames exhibit both spatial (width, height) and temporal (frame index) dimensions. An effective encoding must reflect this inherent structure to facilitate accurate modeling of spatiotemporal dependencies. Recent approaches Wang et al. (2024); Bai et al. (2025), referred to as RoPE-3D, extend RoPE for video structure by splitting the feature channels into three parts to separately encode frame, width, and height positions.
@@
(2) Positional Unbiasedness. A critical yet often overlooked aspect of positional encoding is its impact on attention distribution. As illustrated in Figure 1 (a), RoPE, by design, applies a long-term decay over increasing positional indices, inadvertently introducing a bias that amplifies attention toward later tokens. This issue persists in RoPE-3D, where spatial positions within video frames are unevenly weighted, causing attention to be disproportionately focused on certain areas—typically the bottom-right regions of frames—while suppressing others, which is shown in Figure 1 (a). Such biases distort spatial contextual modeling, leading to suboptimal video comprehension. An effective video positional encoding should mitigate these biases to ensure uniform attention across the entire frame.
@@
(3) Seamless Video-Text Transition. For effective multimodal understanding, an ideal positional encoding should ensure a seamless transition between video and text tokens. However, as demonstrated in Figure 1 (b), RoPE-3D introduces a discontinuity when transitioning from video to text tokens, as the positional indices of text tokens are arbitrarily offset by the maximum position index of the video sequence (determined by the largest of frame count, width, and height, which often vary significantly). This artificial “jump” in the positional encoding space disrupts the smooth flow of information between modalities, hindering the model to establish meaningful cross-modal dependencies.
@@
Based on the above principles, we propose Video Rotary Position Embedding (VRoPE), a novel positional encoding method specifically designed for Video-LLMs. Our approach consists of two key components to satisfy those principles. (1) Symmetric Bias Mitigation: To counteract the attention bias present in RoPE-based encodings, we design a symmetric positional representation that encodes each spatial coordinate from vertices to the center. By distributing attention more uniformly across spatial locations, this method prevents positional distortions and improves overall video understanding. (2) Temporal Centered Arrangement: We propose a center-aligned design that spatially aligns the geometric centers of video frames with the textual arrangement axis, and arranges video frames in temporally ordered progression along the textual positional axis. This transformation not only maintains spatial coherence within video frames but also ensures a smooth transition between video and text tokens, mitigating discontinuities in the positional encoding space.
@@
Overall, VRoPE effectively enhances Video-LLMs by preserving spatiotemporal structure, mitigating attention bias, and ensuring smooth video-text transitions. We conduct extensive experiments on different models and training datasets. Our results demonstrate significant performance improvements over RoPE and RoPE-3D on multiple video benchmarks, covering general video understanding, temporal reasoning, long video comprehension, and video retrieval. These findings establish VRoPE as a robust and efficient positional encoding method tailored for Video-LLMs. We hope this work inspires further research on Video-LLM positional encoding and provides valuable insights for future Video-LLM designs.
@@ 2.1 Video Large Language Models
Recent advancements in Video-LLMs Maaz et al. (2023); Li et al. (2023, 2024b); Jin et al. (2024); Li et al. (2024d); Xu et al. (2024) have significantly enhanced video processing by integrating multiple modalities and employing instruction fine-tuning. Notable innovations include Video-ChatGPT Maaz et al. (2023), which introduced video instruction tuning for text generation, and VideoChat Li et al. (2023) and VideoChat2 Li et al. (2024b), which improved modality alignment via cross-attention and multi-stage bootstrapping etc. Other models, such as Chat-UniVi Jin et al. (2024) and LLaMA-VID Li et al. (2024d), focus on efficient video representations through techniques like token compression and dual-token methods that separate context and content. Additionally, PLLaVA Xu et al. (2024) explores the use of image-pretrained LLaVA models for video tasks, utilizing simple spatial pooling techniques.
@@ 2.2 Multimodal Position Embedding
Most Video-LLMs inherit the default design from LLMs by using Rotary Position Embedding (RoPE) Su et al. (2024) for positional encoding. RoPE encodes relative distance information as absolute position embeddings, offering key advantages like no additional training parameters and improved performance in various tasks Su et al. (2024). It is widely used in modern LLMs due to its ability to extrapolate context length, extending a model’s window size without the need for expensive retraining. However, RoPE’s 1D design, effective for text, overlooks the spatiotemporal structure of video data, limiting its suitability for Video-LLMs. To address this, several approaches have adapted RoPE for video. For instance, RoPE-2D Agrawal et al. (2024); Wang et al. (2024) extends the encoding to capture spatial relationships in video frames, while RoPE-3D Wang et al. (2024); Bai et al. (2025) divides the channel dimension into three groups to better represent the spatiotemporal dimensions.
@@ 3.1 Preliminary: Rotary Position Embedding
where base is a hyperparameter, $d$ is the feature dimension, and $j=[0,1,...,d/2-1]$ denotes the index of each feature channel.
@@ 3.2 RoPE for Video-LLMs
In Video-LLMs, video frames are typically processed by vision encoders (e.g., ViTs Alexey (2020) or CNNs He et al. (2016)) and transformed into a sequence of visual tokens. These visual tokens are then concatenated with text tokens and fed into an LLM backbone.
In most existing approaches, video tokens are treated as a simple 1D sequence, with position indices assigned in an increasing order, similar to text. However, this naive approach, referred to as RoPE, overlooks the inherent spatiotemporal structure of video data. Flattening video frames this way disrupts spatiotemporal structure and leads to inefficient position usage. Unlike text, video tokens carry less dense semantic information, and their excessive sequence length can weaken contextual dependencies, making long-range understanding harder.
@@ 3.3 RoPE-3D for Video-LLMs
[…]
@@
[…]
@@ 3.4 Problem Analysis
Table 1: Average attention weights at the video-text boundary on Video-MME. We use the subsequent text instruction as the query and video/text tokens as keys. Note that text-to-video attention weights of RoPE-3D are an order of magnitude lower than other methods, indicating its positional discontinuity between video and text.
Method Text-to-Text Text-to-Video
RoPE 1.41e-2 2.08e-4
RoPE-3D 1.27e-2 5.12e-5
VRoPE (Ours) 1.32e-2 3.70e-4
@@
Figure 3: Left: the overall architecture of a typical Video-LLM. In this work, our improvements primarily target the positional embedding component of the LLM to enhance its video understanding capability. Right: method illustration of VRoPE. (a) We first apply symmetric arrangement to mitigate positional bias in video frames. The RoPE frequencies are uniformly allocated to the four dimensions. (b) We propose to use temporal centered arrangement in video frames to form a seamless video-text transition, which enables video input of arbitrary length without causing discontinuity.
@@
RoPE-3D introduces separate positional encodings for spatial (width, height) and temporal (frame index) dimensions. However, when video tokens are concatenated with subsequent text tokens, their positional indices do not follow a smooth transition. Instead, text tokens inherit positional indices that are arbitrarily offset by the maximum position value across spatial $(W,H)$ and temporal dimensions $T$, i.e., $\max(W,H,T)$. This results in an artificial “jump” in the positional encoding space when transitioning from video to text tokens. The discontinuity creates an abrupt and non-uniform gap between the final video token and the subsequent text token. As is shown in Table 1, text-to-video attention weights of RoPE-3D at the video-text boundary are an order of magnitude lower than RoPE and VRoPE, which demonstrates that the discontinuity in position embedding will affect the attention weights. Further, the magnitude of this gap depends on video dimensions rather than being a fixed offset, making it inconsistent across different video-text samples. Such a discrepancy can degrade the model’s ability to establish seamless contextual dependencies across modalities. This issue is particularly problematic in long videos, as the increasing frame count $T$ exacerbates the positional gap, which will be further discussed in Section 5.3.
@@ 4 Method: VRoPE
In this section, we introduce Video Rotary Position Embedding (VRoPE), a novel positional encoding method tailored for Video-LLMs. Our approach addresses the inherent limitations of RoPE-3D, including positional attention bias and cross-modal positional discontinuity, by leveraging a combination of Symmetric Bias Mitigation and Temporal Centered Arrangement. The overall framework of VRoPE is illustrated in Figure 3.
@@ 4.1 Symmetric Bias Mitigation
As discussed in Section 3.4, both RoPE and RoPE-3D employ a single positional arrangement direction when encoding features within video frames (e.g., row-major scanning for RoPE and top-left to bottom-right ordering for RoPE-3D), inevitably introducing positional attention bias. To address this limitation, we propose Symmetric Bias Mitigation as illustrated in Figure 3 (a).
Specifically, we design a unified symmetric positional arrangement paradigm applicable to arbitrary dimensions. For textual tokens represented as points, their inherent symmetry is preserved. For one-dimensional sequences, we adopt bidirectional positional indexing starting from both endpoints (similar to bidirectional modeling in language models). For two-dimensional planes (i.e., video frames), we implement a four-directional symmetric arrangement extending from frame vertices toward the center. This scheme naturally extends to three-dimensional space with eight-vertex symmetry, etc. Given an input video frame of size $(W,H)$, we compute four symmetric directional positional arrangements as follows:
$\begin{pmatrix}u_{1}\\ u_{2}\\ u_{3}\\ u_{4}\end{pmatrix}=\begin{pmatrix}w+h\\ w-h\\ -w-h\\ -w+h\end{pmatrix}.$ (6)
Considering that RoPE employs different frequencies across channels, we strategically allocate frequencies to these four symmetric positional indices in a uniform manner. This design enables distinct positional arrangement directions to model features through different RoPE frequencies (high, medium and low).
4.2 Temporal Centered Arrangement
While Symmetric Bias Mitigation effectively alleviates positional bias, the inherent discontinuity between video and textual modalities persists. To address this challenge, we propose the Temporal Centered Arrangement for positioning video frames. Given that textual positions inherently satisfy $u_{1}=u_{2}=u_{3}=u_{4}$ (demonstrating isotropic symmetry), we first align the geometric center of each video frame with the textual arrangement axis through coordinate transformation. Specifically, for a video of size $(W,H,T)$ with an initial position index $p_{start}$ (i.e., the last position id + 1 of the previous text), this process can be denoted as:
$\begin{pmatrix}v_{1}\\ v_{2}\\ v_{3}\\ v_{4}\end{pmatrix}=\begin{pmatrix}u_{1}\\ u_{2}+H-1\\ u_{3}+H+W-2\\ u_{4}+W-1\end{pmatrix}+p_{start}.$ (7)
Subsequently, we systematically arrange frame positions along the temporal dimension using the following formulation:
$v_{j}^{t}=v_{j}+t(H+W-1),$ (8)
where $t$ is the frame index. This configuration ensures that: (1) The central position of each video frame coincides with the textual arrangement axis, and (2) Sequential frames naturally extend along the textual positional progression direction through temporal ordering. Consequently, the temporal expansion axis of video sequences becomes intrinsically aligned with the positional growth direction of text tokens, which means that arbitrary length of video input does not affect the continuity between video and text.
@@ 4.2 Temporal Centered Arrangement
$\text{VRoPE}_{j}($ $\mathbf{x},v_{1}^{t},v_{2}^{t},v_{3}^{t},v_{4}^{t})$ (9)
=$&\text{RoPE}_{j}(\mathbf{x},v_{1}^{t})_{j=4k}\\ &\text{RoPE}_{j}(\mathbf{x},v_{2}^{t})_{j=4k+1}\\ &\text{RoPE}_{j}(\mathbf{x},v_{3}^{t})_{j=4k+2}\\ &\text{RoPE}_{j}(\mathbf{x},v_{4}^{t})_{j=4k+3}\end{aligned}\right.$
where $k\in\left\{0,1,2,...\right\}$. For text tokens, we retain the original RoPE encoding structure (Eq. 5) to ensure compatibility with LLMs. Further discussions can be found in Appendix A.

### v3

@@ Abstract
Rotary Position Embedding (RoPE) has shown strong performance in text-based Large Language Models (LLMs), but extending it to video remains a challenge due to the intricate spatiotemporal structure of video frames. Existing adaptations, such as RoPE-3D, attempt to encode spatial and temporal dimensions separately but suffer from two major limitations: positional bias in attention distribution and disruptions in video-text transitions. To overcome these issues, we propose Video Rotary Position Embedding (VRoPE), a novel positional encoding method tailored for Video-LLMs. Our approach restructures positional indices to preserve spatial coherence and ensure a smooth transition between video and text tokens. Additionally, we introduce a more balanced encoding strategy that mitigates attention biases, ensuring a more uniform distribution of spatial focus. Extensive experiments on Vicuna and Qwen2 across different model scales demonstrate that VRoPE consistently outperforms previous RoPE variants, achieving significant improvements in video understanding, temporal reasoning, and retrieval tasks. Code will be available at https://github.com/johncaged/VRoPE
@@ 1 Introduction
To optimally encode positional relationships in Video-LLMs, we identify three key properties that an ideal video positional encoding method should satisfy:
@@
(1) Spatiotemporal Structure Modeling. Unlike text, where positional relationships are strictly one-dimensional, video frames exhibit both spatial (width, height) and temporal (frame index) dimensions. An effective encoding must reflect this inherent structure to facilitate accurate modeling of spatiotemporal dependencies. Recent approaches Wang et al. (2024) have attempted to extend RoPE for video structure. These methods, which we refer to as RoPE-3D, split the feature channels into 3 parts to separately encode frame, width, and height positions.
@@
(2) Positional Unbiasedness. A critical yet often overlooked aspect of positional encoding is its impact on attention distribution. As illustrated in Figure 1 (a), RoPE, by design, applies a long-term decay over increasing positional indices, inadvertently introducing a bias that amplifies attention toward later tokens. This issue persists in RoPE-3D, where spatial positions within video frames are unevenly weighted, causing attention to be disproportionately focused on certain areas—typically the bottom-right regions of frames—while suppressing others, which is shown in Figure 1 (a). Such biases distort spatial contextual modeling and can lead to suboptimal video comprehension. An effective video positional encoding should mitigate these biases to ensure uniform attention across the entire frame.
@@
(3) Seamless Video-Text Transition. For effective multimodal understanding, an ideal positional encoding should ensure a seamless transition between video and text tokens. However, as demonstrated in Figure 1 (b), RoPE-3D introduces a discontinuity when transitioning from video to text tokens, as the positional indices of text tokens are arbitrarily offset by the maximum position index of the video sequence (determined by the largest of frame count, width, and height, which often vary significantly). This artificial “jump” in the positional encoding space disrupts the smooth flow of information between modalities, making it harder for the model to establish meaningful cross-modal dependencies.
@@
Based on the above principles, we propose Video Rotary Position Embedding (VRoPE), a novel positional encoding method specifically designed for Video-LLMs. Our approach consists of two key components to satisfy those principles. (1) Cross-Modal Continuity Rotation: We introduce a spatial rotation transformation that redistributes RoPE-3D’s positional indices while preserving local spatial structures. This transformation not only maintains spatial coherence within video frames but also ensures a smooth transition between video and text tokens, mitigating discontinuities in the positional encoding space. (2) Symmetric Bias Mitigation: To counteract the attention bias present in RoPE-based encodings, we design a symmetric positional representation that encodes each spatial coordinate with both positive and negative components. By distributing attention more uniformly across spatial locations, this method prevents positional distortions and improves overall video understanding.
@@
Overall, VRoPE effectively enhances Video-LLMs by preserving spatiotemporal structure, mitigating attention bias, and ensuring smooth video-text transitions. We conduct extensive experiments on Vicuna Chiang et al. (2023) and Qwen2 Yang et al. (2024) across different model scales (1.5B and 7B parameters). Our results demonstrate significant performance improvements over RoPE, RoPE-3D, and other positional encoding variants on multiple video benchmarks, covering general video understanding, temporal reasoning, long video comprehension, and video retrieval. These findings establish VRoPE as a robust and efficient positional encoding method tailored for Video-LLMs.
@@ 2.1 Video Large Language Models
Recent advancements in Video-LLMs have significantly enhanced video processing by integrating multiple modalities and employing instruction fine-tuning. Notable innovations include Video-ChatGPT Maaz et al. (2023), which introduced video instruction tuning for text generation, and VideoChat Li et al. (2023) and VideoChat2 Li et al. (2024b), which improved modality alignment via cross-attention and multi-stage bootstrapping etc. Other models, such as Chat-UniVi Jin et al. (2024) and LLaMA-VID Li et al. (2024d), focus on efficient video representations through techniques like token compression and dual-token methods that separate context and content. Additionally, PLLaVA Xu et al. (2024) explores the use of image-pretrained LLaVA models for video tasks, utilizing simple spatial pooling techniques.
@@ 2.2 Multimodal Position Embedding
Most Video-LLMs inherit the default design from LLMs by using Rotary Position Embedding (RoPE) Su et al. (2024) for positional encoding. RoPE encodes relative distance information as absolute position embeddings, offering key advantages like no additional training parameters and improved performance in various tasks Su et al. (2024). It is widely used in modern LLMs due to its ability to extrapolate context length, extending a model’s window size without the need for expensive retraining. However, RoPE’s 1D design, effective for text, overlooks the spatiotemporal structure of video data, limiting its suitability for Video-LLMs. To address this, several approaches have adapted RoPE for video Su (2024); Wang et al. (2024). For instance, RoPE-2D Agrawal et al. (2024); Wang et al. (2024) extends the encoding to capture spatial relationships in video frames, while RoPE-3D Wang et al. (2024) divides the channel dimension into three groups to better represent the spatiotemporal dimensions.
@@ 3.1 Preliminary: Rotary Position Embedding
where base is a hyperparameter, $d$ is the feature dimension, and j=[0,1,,d/21]j=[0,1,...,d/2-1] denotes the index of each feature channel.
Figure 2: Overall architecture of a typical Video-LLM. The input videos are first encoded by vision encoders, followed by a connector layer. The resulting video features are then flattened and concatenated with text embedding tokens to form a multimodal input sequence for the LLM. In this work, our improvements primarily target the positional embedding component of the LLM to enhance its video understanding capability.
@@ 3.2 RoPE for Video-LLMs
In Video-LLMs, video frames are typically processed by vision encoders (e.g., ViTs Alexey (2020) or CNNs He et al. (2016)) and transformed into a sequence of visual tokens. These visual tokens are then concatenated with text tokens and fed into an LLM backbone. The overall architecture of typical Video-LLMs is denoted in Figure 2.
In most existing approaches, video tokens are treated as a simple 1D sequence, with position indices assigned in an increasing order, similar to text. However, this naive approach, referred to as RoPE, overlooks the inherent spatiotemporal structure of video data. Flattening video frames this way disrupts spatiotemporal structure and leads to inefficient token usage. Unlike text, video tokens carry less dense semantic information, and their excessive sequence length can weaken contextual dependencies, making long-range understanding harder.
@@ 3.3 RoPE-3D for Video-LLMs
$\text{RoPE-3D}_{j}(\mathbf{x},w,h,t)=\left\{\begin{aligned} &\text{RoPE}_{j}(\mathbf{x},w),j\in D_{w}\\ &\text{RoPE}_{j}(\mathbf{x},h),j\in D_{h}\\ &\text{RoPE}_{j}(\mathbf{x},t),j\in D_{t}\end{aligned}\right$. (4)
@@
$\text{RoPE-3D}_{j}(\mathbf{x},m,m,m)\equiv\text{RoPE}_{j}(\mathbf{x},m)$ (5)
@@ 3.4 Problem Analysis
RoPE-3D introduces separate positional encodings for spatial (width, height) and temporal (frame index) dimensions. However, when video tokens are concatenated with subsequent text tokens, their positional indices do not follow a smooth transition. Instead, text tokens inherit positional indices that are arbitrarily offset by the maximum position value across spatial $(W,H)$ and temporal dimensions $T$, i.e., $\max(W,H,T)$. This results in an artificial “jump” in the positional encoding space when transitioning from video to text tokens. The discontinuity creates an abrupt and non-uniform gap between the final video token and the subsequent text token. The magnitude of this gap depends on video dimensions rather than being a fixed offset, making it inconsistent across different video-text samples. Such a discrepancy can degrade the model’s ability to establish seamless contextual dependencies across modalities. This issue is particularly problematic in long videos, as the increasing frame count $T$ exacerbates the positional gap.
@@ 4 Method: VRoPE
In this section, we introduce Video Rotary Position Embedding (VRoPE), a novel positional encoding method tailored for Video-LLMs. Our approach addresses the inherent limitations of RoPE-3D, including spatiotemporal attention bias and cross-modal positional discontinuity, by leveraging a combination of Cross-Modal Continuity Rotation and Symmetric Bias Mitigation. The overall framework of VRoPE is illustrated in Figure 4.
Figure 4: Method illustration of VRoPE.
4.1 Cross-Modal Continuity Rotation
To mitigate the cross-modal positional discontinuity introduced when concatenating video and text tokens, we propose a spatial rotation transformation that redistributes positional indices while preserving the inherent spatial structure. Specifically, we transform the width and height coordinates into a rotated space as follows:
$\begin{pmatrix}u\\ v\end{pmatrix}=\begin{pmatrix}w+h\\ w-h+H-1\end{pmatrix}$ (6)
where the term H1H-1 ensures that $v$ remains non-negative. As is shown in Figure 4 (a), this transformation provides two key benefits: (1) it preserves the local spatial relationships of tokens within the frame, and (2) it ensures a smooth transition from video to text tokens by aligning their positional indices along diagonal axes in the $(u,v)$ space. Consequently, the model naturally reduces abrupt jumps in positional encoding, maintaining consistent contextual dependencies across modalities.
@@ 4.2 Symmetric Bias Mitigation
While Cross-Modal Continuity Rotation alleviates discontinuity, the issue of spatial attention bias persists. To address this, we introduce Symmetric Bias Mitigation, which ensures a more uniform attention distribution across spatial dimensions.
As is shown in Figure 4 (b), instead of using scalar positional values, we represent each spatial position $u,v$ as a symmetric vector to capture bidirectional positional dependencies, i.e., $u=(u_{+},u_{-})$ and $v=(v_{+},v_{-})$. To achieve this, we evenly partition the feature dimensions into four equal parts, assigning each part to a different positional encoding term. For a video of size $(W,H,T)$ with an initial position index $p_{start}$, its first frame’s position indices are computed as follows:
$\begin{pmatrix}u_{+}\\ u_{-}\\ v_{+}\\ v_{-}\end{pmatrix}=\begin{pmatrix}u\\ -u\\ v\\ -v\end{pmatrix}+\begin{pmatrix}0\\ bias\\ 0\\ bias\end{pmatrix}+p_{start}$ (7)
where bias=H+W2bias=H+W-2 ensures non-negative values. After processing the first frame, the position index $p_{start}$ is updated to $p_{start}+H+W-1$, and subsequent frames follow the same encoding scheme.
This encoding strategy ensures that the decay pattern is counterbalanced across positive and negative position pairs. Consequently, the distance between frame tokens at different spatial positions—and between frame tokens and subsequent text tokens—remains more consistent. As is shown in Figure 3 (c), this alleviates the skewed focus on specific regions and improves overall video understanding.
@@
$\text{VRoPE}_{j}($ $\mathbf{x},u_{+},u_{-},v_{+},v_{-})$ (8)
$=\left\{\begin{aligned} &\text{RoPE}_{j}(\mathbf{x},u_{+})_{j=4k}\\ &\text{RoPE}_{j}(\mathbf{x},u_{-})_{j=4k+1}\\ &\text{RoPE}_{j}(\mathbf{x},v_{+})_{j=4k+2}\\ &\text{RoPE}_{j}(\mathbf{x},v_{-})_{j=4k+3}\end{aligned}\right$.
[…]
4.3 Discussion
In addition to preserving the video structure, ensuring a continuous transition between video and text tokens, and mitigating spatial attention bias, VRoPE also exhibits the following characteristics.
Implicit Temporal Modeling.
While VRoPE does not directly encode frame indices, it implicitly captures temporal cues by arranging frame tokens in an ascending sequence and applying symmetric positional transformations. The temporal correlations among the frames are implicitly captured through the opposite direction design, ensuring the model differentiate temporal progressions.
Dimensional Adaptability.
A key advantage of VRoPE is its ability to degenerate into lower-dimensional embeddings without altering its fundamental structure. Unlike methods that assign separate feature channels for each coordinate, VRoPE employs linear combinations of the original coordinates, allowing any dimension set to 1 to seamlessly adapt into lower-dimension form. For instance, when $H=1$, the encoded positions simplify to (w,w,w,w)(w,-$w,$-w), effectively reducing to a 1D form—unlike previous methods that rely on separate encodings, such as $(w,0)$. This property is particularly beneficial for adapting pre-trained model’s positional encodings from images (2D) or videos (3D) to data of varying dimensions without disrupting the original encoding scheme. Consequently, models can transfer more effectively across modalities while preserving consistent positional behavior.